### v4

# Global Compression Commander: Plug-and-Play Inference Acceleration for High-Resolution Large Vision-Language Models

Abstract
[…]
Large vision-language models (LVLMs) excel at visual understanding and reasoning, but face efficiency challenges due to quadratic complexity in processing long multimodal contexts. While token compression techniques can reduce computational costs, existing approaches are designed for single-view LVLMs and fail to consider the unique multi-view characteristics of recent high-resolution LVLMs with dynamic tiling. While existing methods treat all tokens uniformly, our analysis reveals that global thumbnails can naturally guide the compression of local crops by providing holistic context for informativeness evaluation. In this paper, we first analyze dynamic tiling strategy comprehensively, revealing both the complementary nature between thumbnails and crops, and the distinctive characteristics across different crops. Based on our observations, we propose “Global Compression Commander” (i.e., GlobalCom2), a novel plug-and-play token compression framework for HR-LVLMs. GlobalCom2 leverages thumbnail as the “commander” to guide the compression process of local crops, adaptively preserving informative details while eliminating redundancy. Extensive experiments show that GlobalCom2 maintains over 90% performance while compressing 90% visual tokens, reducing FLOPs and peak memory to 9.1% and 60% respectively across multiple benchmarks. Our code is available at https://github.com/xuyang-liu16/GlobalCom2.
Figure 1: Design philosophy of “global-to-local” guided token compression. GlobalCom2 evaluates the information richness of crops from a global perspective to preserve informative regions while removing redundancy.
1 Introduction
By bridging visual encoders with Large Language Models (LLMs) [40, 35, 1], Large Vision-Language Models (LVLMs) [27, 28, 2, 24, 51] have achieved remarkable progress in multiple tasks. However, the large number of visual tokens leads to prohibitively expensive computation during inference [8, 38, 18, 13], especially for high-resolution images, limiting LVLMs’ practical applications.
To tackle the challenge of inference speed in LVLMs, researchers have primarily focused on model-centric approaches like knowledge distillation [4] and quantization [43], which typically require model retraining, consuming substantial computational resources thus limiting their practical applicability. Recent efforts have shifted towards data-centric approaches, exemplified by token compression methods [5, 47, 38, 31, 16], which consider the redundancy in LVLMs’ input visual signals to enhance inference speed by reducing processed tokens while preserving essential information. Most of these methods are architecture-agnostic and training-free, achieving optimal efficiency-accuracy trade-off while being complementary to model-centric approaches [8, 53, 52].
However, current token compression methods were designed for LVLMs processing single-view images (entire image without crop decomposition, as shown in Figure 1 top). With growing demand for high-resolution image understanding, advanced HR-LVLMs like LLaVA-NeXT [29] and InternVL 1.5 [9] adopt dynamic tiling, which divides an image into several crops encoded individually and concatenated with tokens from a global thumbnail (in Figure 1 top). While this multi-view approach enables more comprehensive visual understanding, it further increases the number of visual tokens and computational overhead.
Directly applying existing token compression methods to all views might seem straightforward, but such uniform compression overlooks two critical structural properties revealed by our analysis: (1) The inherent role of global thumbnails in extracting holistic context and guiding general vision understanding, demonstrated in Figure 2, and (2) The information richness disparity across crops quantified in Figure 3 (e.g., 2.4% average performance gap between discarding most/least informative crops). These oversights lead to over-compression of information-rich regions and disrupted visual-semantic hierarchy (see Figure 6).
To establish theoretically grounded compression principles for dynamic tiling-based HR-LVLMs, we first conduct systematic analysis to investigate the roles of global thumbnails and local crops in HR-LVLM image understanding capabilities. Through extensive empirical studies, we identify two key observations for token compression: Visual tokens from thumbnails and crops serve complementary roles - thumbnail tokens capture holistic context while local crop tokens offer fine-grained details, enabling thumbnails to guide holistic token importance evaluation from a global perspective. Through global context, visual tokens from different crops exhibit varying degrees of informativeness and visual redundancy, necessitating differentiated compression intensity to minimize information loss.
Figure 2: Effects of global thumbnail and local crops. Performance (%) denotes relative scores vs. vanilla LLaVA-NeXT-7B.
Building on these observations, we propose “Global Compression Commander” (i.e., GlobalCom2), which follows a “global-to-local” guided token compression design philosophy, tailored for HR-LVLMs with dynamic tiling strategy. As shown in Figure 1, GlobalCom2 first leverages holistic information from global thumbnail perspective to evaluate the information richness of each crop, adaptively allocating compression budgets across different crops. It then performs token compression for local crops through comprehensive token evaluation from both global and local perspectives, achieving differentiated compression degrees across regions. Through this “global-to-local” design, GlobalCom2 achieves training-free adaptive token compression, which selectively preserves significant visual information while removing redundancy at both global and local perspectives for HR-LVLMs with dynamic tiling.
To summarize, our main contributions are three-fold:
  • We empirically analyze the hierarchical nature of dynamic tiling strategy by demonstrating the distinct and complementary roles between thumbnails and crops, as well as the inherent characteristics among local crops, leading to two key observations for token compression.
  • We propose a “global-to-local” guided token compression framework GlobalCom2, which adaptively adjusts compression intensity based on crops’ information richness and preserves semantically important tokens with fine-grained details across the entire image.
  • Extensive experiments on LLaVA-NeXT and LLaVA-OneVision demonstrate superior efficiency-performance trade-offs, maintaining 90% performance with 90% token reduction, while achieving substantial memory savings and throughput improvement.
Figure 3: Effects of different local crops. Crop importance is quantified through accumulating attention scores between thumbnail patches and the [CLS] token within each crop region.
2 Related Work
2.1 High-resolution LVLMs
LVLMs generally combine a vision encoder and a large language model (LLM) decoder, bridged by a visual projector that maps visual features into the LLM’s embedding space. Representative works include BLIP-2 [22] which utilizes a Q-Former as the visual projector, and LLaVA-1.5 [28] which employs a multi-layer perceptron (MLP) projector to integrate visual and textual information.
Previous LVLMs [28, 2] resize images to a fixed resolution before processing, which often leads to shape distortion and loss of local details in high-resolution images. To address this, a series of LVLMs have been proposed to handle high-resolution images, which can be categorized into three groups: (1) Hybrid resolution methods using dual visual encoders, represented by Mini-Gemini [24] and LLaVA-HR [34]. (2) Native resolution methods adopting NaViT-style [11] encoders, represented by Qwen2-VL [42]. (3) Dynamic tiling methods, represented by LLaVA-NeXT [29] and InternVL-1.5 [9], which split images into regions and process them with a single encoder. Due to its efficiency and effectiveness in vision encoding, dynamic tiling has been widely adopted in subsequent works [30, 6, 21].
While increasing local visual tokens enhances detail recognition and hallucination suppression, it also introduces computational challenges in inference speed and memory usage. In this work, we focus on improving the efficiency of dynamic tiling-based HR-LVLMs.
2.2 Token Compression for LVLMs
Token compression, which aims to reduce the sequence length of tokens for computation efficiency, has been widely adopted for model acceleration [37, 25, 3]. For LVLMs, recent works have focused on training-free token compression paradigms, which can be broadly categorized into two streams: (1) Token compression during Vision Encoding stage [38, 45, 16]: For example, FasterVLM [52] utilizes [CLS] attention scores from the visual encoder to re-rank and retain top visual tokens. (2) Token compression during LLM Decoding stage [8, 46, 26]: Representative works include FastV [8], which prunes tokens based on self-attention scores in LLM, and SparseVLM [53], which adaptively prunes visual tokens based on their attention scores with text tokens. However, when applying these methods to HR-LVLMs with dynamic tiling, they treat all visual tokens with uniform compression strategies, overlooking the distinct roles of global thumbnail and local crop tokens in image understanding for HR-LVLMs.
In this paper, unlike existing token compression methods, we propose a “global-to-local” guided compression strategy that leverages global thumbnail as a “commander” to guide token compression in HR-LVLMs.
3 Analysis of HR-LVLM with Dynamic Tiling
3.1 Preliminary
We take LLaVA-NeXT [29], a widely-adopted HR-LVLM with dynamic tiling, as an example to conduct our analysis.
Model Architecture. Given the input image and textual instructions, LLaVA-NeXT generates instruction-following responses in an autoregressive manner through two stages:
  • Vision Encoding: LLaVA-NeXT adopts CLIP-ViT [36] to transform raw pixels into unified visual representations, followed by a two-layer MLP Projector that aligns visual embeddings with the language embedding space.
  • LLM Decoding: The decoder-only LLM takes concatenated image and prompt tokens as input and decodes them through next-token prediction to generate responses.
Dynamic Tiling Input Pre-processing. LLaVA-NeXT processes arbitrary-resolution images through adaptive dynamic cropping using seven grid templates: $\{2\times 2,1\times\{2,3,4\},\{2,3,4\}\times 1\}$. For an input image of size $W\times H$, it is scaled to $(336\times a)\times(336\times b)$ based on the selected template, producing $n=a\times b$ local crops $\mathbf{X}^{L}$ of size $336\times 336$ (with padding if necessary) and a thumbnail $\mathbf{X}^{G}$. With $N$ tokens per crop, the total token length is $(1+n)\times N$.
Visual Tokens Post-processing. Following [28], LLaVA-NeXT removes padding tokens and uses special tokens to mark image boundaries, preserving aspect ratio information while improving efficiency.
3.2 Stepping into Dynamic Tiling Mechanism
HR-LVLMs process high-resolution images using dynamic tiling with thumbnails and crops. We analyze their characteristics to guide token compression design.
(I) Functions of Thumbnail and Crops. We begin with an exploratory experiment to investigate how thumbnail and crops contribute to image understanding in HR-LVLMs by using them separately as input to LLaVA-NeXT-7B model.
As shown in Figure 2, using global thumbnail alone yields superior performance on general visual perception benchmarks such as SQA [33], MME [12], MMBench [32], and MM-Vet [48]. This advantage stems from global thumbnails providing holistic visual information through complete ViT encoding. In contrast, using only local crops, where ViT independently encodes each crop, shows inferior performance on these general visual perception tasks. However, local crops excel in fine-grained perception tasks like TextVQA [39] (VQAT, testing text-centric visual understanding) and POPE [23] (testing model hallucination) by providing detailed features. Therefore, we identify that:
Observation ❶: Global thumbnails and local crops serve complementary functions in HR-LVLMs with dynamic tiling: the former acts as a “comprehensive visual extractor” for holistic representations, while the latter serves as a “detailed visual capturer” for fine-grained representations.
Figure 4: Overall framework of GlobalCom2. GlobalCom2 guides token compression for HR-LVLMs through: 1) compressing thumbnail tokens (blue path); and 2) compressing crop tokens (yellow paths) by: (a) adaptively adjusting compression intensity based on visual information richness at global view, and (b) performing compression according to token informativeness at global and local views.
(II) Information Richness in Different Crops. Given that dynamic tiling leads to distinct visual representations across crops, we investigate the variance through both qualitative and quantitative perspectives.
Through visual inspection of input image in Figure 4, we observe that the upper two crops contain significant visual content (e.g., football players), while the lower two crops exhibit considerable visual redundancy (predominantly grass areas). To quantitatively validate this observed variation in informational content, we leverage insights from previous works [25, 16], which demonstrated that attention scores between [CLS] and patch tokens effectively reflect token importance. As shown in the upper middle of Figure 4, our visualization of attention scores between patch tokens and the [CLS] token from CLIP-ViT’s last layer reveals notably higher attention values in the upper regions, confirming richer semantic information in upper crops from a global perspective.
Building upon this qualitative analysis, we conduct quantitative studies to validate our observations. Since attention scores with the [CLS] token effectively measure token-level semantic richness, we leverage these scores to assess crop importance by by computing the sum of partitioned attention scores within each crop region. To examine crop contributions in HR-LVLMs, we conduct experiments on LLaVA-NeXT-7B by selectively dropping local crops (i.e., dropping the most/least important one). As shown in Figure 3, dropping the most versus least important crops leads to significant performance gaps across visual understanding tasks, with an average drop of 2.4% across six benchmarks and particularly notable in VQAT (5.2% gap). Based on the above analysis, we identify that:
Observation ❷: Local crops exhibit varying information richness, leading to different contributions to the overall visual understanding of HR-LVLMs with dynamic tiling, with visually informative crops being particularly crucial for capturing fine-grained local details.
4 Global Compression Commander
As discussed in Section 3, HR-LVLMs’ dynamic tiling strategy leads to significant token length increase ( in LLaVA-NeXT [29], 10× in LLaVA-OneVision [21]), where the quadratic complexity of self-attention [41] becomes a major bottleneck. To address this challenge, we aim to directly improve HR-LVLMs’ inference speed through visual token compression, as illustrated in Figure 4.
Motivated by observation in Section 3.2 that global thumbnail serves as a “comprehensive visual extractor” providing visual representations from a holistic perspective, we leverage it as a “Global Compression Commander” (GlobalCom2) to guide token compression for HR-LVLMs, as illustrated in Figure 4. In what follows, we first elaborate on how our GlobalCom2 guides token compression for the global thumbnail.
Since visual token compression aims to preserve informative tokens while eliminating redundant representations in visual content, the key for global thumbnails lies in identifying tokens that capture essential holistic information. Given that the [CLS] token serves as an effective global image representation [25, 16], GlobalCom2 leverages the last ViT layer’s attention map to compute average attention between each global thumbnail token and the [CLS] token (blue path in Figure 4). Specifically, for 1D token sequence $\mathbf{X}^{G}$ of length $N$ in the global thumbnail, the importance score ${s}_{i}^{G}$ for the $i$-th token is:
$\mathbf{s}_{i}^{G}=\frac{\exp{(\mathbf{q}^{\texttt{CLS}}{\mathbf{K}}_{i}^{\top}/\sqrt{D})}}{\sum^{N}_{i=1}\exp{(\mathbf{q}^{\texttt{CLS}}{\mathbf{K}}_{i}^{\top}/\sqrt{D})}},$ (1)
where $\mathbf{q}^{\texttt{CLS}}$ and $\mathbf{K}\in\mathbb{R}^{N\times D}$ are query of [CLS] and keys of $\mathbf{X}^{G}$. Based on these scores, tokens with higher ${s}^{G}$ capture richer semantic information and should be preserved. Given a preset token retention ratio $R$ (%), GlobalCom2 preserves the top-$k$ ($k=R\times N$) tokens ranked by ${s}^{G}$:
$\mathbf{X}^{G}\rightarrow\mathbf{\hat{X}}^{G}=\text{TopK}(\mathbf{X}^{G},s^{G},R\times N)$ (2)
where $\mathbf{\hat{X}}^{G}$ is the compressed thumbnail sequence. This training-free compression mechanism reduces computational cost by preserving only the informative tokens.
Due to dynamic tiling’s multi-partitioning process, local crops produce multiple independent token sequences containing distinct visual contents, with each crop exhibiting different levels of redundancy and collectively leading to longer sequences compared to the global thumbnail. Inspired by observation , local crops with different semantic richness warrant different token compression degrees. Specifically, semantically rich crops should retain more tokens to preserve crucial local details, while less informative ones can be more aggressively compressed to help LLMs focus on informative visual signals.
To implement such adaptive compression, motivated by observations -, we aim to guide the token compression of local crops from a global perspective. Specifically, GlobalCom2 employs two aspects of global guidance: (a) Adaptive Compression Adjustment that dynamically adapts compression intensity for each crop based on its information richness, and (b) Holistic Token Evaluation that assesses token informativeness from both local and global perspectives. We detail these two global guidance below.
4.1 Adaptive Compression Adjustment
As presented in the bottom-left of Figure 4, GlobalCom2 first analyzes the semantic contribution of each local crop to the overall visual information, and accordingly applies adaptive token compression at different levels.
To achieve differentiated token compression for local crops, GlobalCom2 adaptively adjusts the compression degrees for each crop based on its informativeness. Specifically, for the $j$-th crop, we define its information richness score $s_{j}^{G}$ from the global perspective by accumulating patch-to-[CLS] attention scores over its corresponding region in the global thumbnail: $s_{j}^{G}=\sum_{i\in\text{crop}_{j}}s^{G}_{i}$, where $s^{G}_{i}$ denotes the informativeness score of each token in the crop region from the thumbnail. To stabilize $s_{j}^{G}$, we compute $\tilde{s}_{j}=(s_{j}^{G}-\max(s_{j}^{G}))/\tau$ with a temperature parameter $\tau=10$, and then compute the relative importance weight $\sigma_{j}$ of the $j$-th crop via softmax across all $n$ crops:
$\sigma_{j}=\frac{\exp(\tilde{s}_{j})}{\sum_{l=1}^{n}\exp(\tilde{s}_{l})+\epsilon},$ (3)
where $\epsilon=10^{-8}$ prevents division by zero. Finally, we adjust the preset retention ratio $R(\%)$ according to the computed importance weight $\sigma_{j}$. This adjustment ensures crops with higher importance preserve more tokens while maintaining the average retention ratio:
$r_{j}=R\times\left(1+\sigma_{j}-\frac{1}{n}\right),$ (4)
where $\sigma_{j}-\frac{1}{n}$ represents the relative deviation from average importance, leading to more aggressive compression for less important regions (values exceeding 100% are capped).
To this end, GlobalCom2 optimally allocates compression degrees (i.e., $\{r_{1},r_{2},r_{3},r_{4}\}$ in Figure 4) based on each crop’s information richness from global view, enabling differentiated token compression degrees across crops.
4.2 Holistic Token Evaluation
After determining the optimal compression budgets, GlobalCom2 evaluates the importance of tokens within each crop and preserves the essential ones, as shown in the bottom-right of Figure 4.
After independent ViT encoding of each local crop, the attention values between patch tokens and the [CLS] token in its corresponding attention map reflect each token’s semantic importance within the crop. Similar to global analysis, GlobalCom2 computes a set of local importance scores $\{s_{j}^{L}\}_{j=1}^{n}$, where $s_{j,i}^{L}$ denotes the score of the $i$-th token in the $j$-th crop based on its own final-layer attention map. Here, each $s_{j}^{L}$ represents the attention scores after excluding padding tokens. However, since local crops are encoded independently during vision encoding stage, these local scores only reflect token importance within individual crops, failing to capture their significance in the full image context. This local-only evaluation poses a potential challenge, as semantically crucial elements that bridge different image regions might be overlooked.
Motivated by observation that global thumbnails provide holistic visual representations, GlobalCom2 extends local token evaluation with global context to achieve more comprehensive importance assessment. As shown in Figure 4, GlobalCom2 reshapes the 1D attention scores ${s}^{G}$ of $\mathbf{X}^{G}$ into 2D format and bilinearly interpolates them to match the original image dimensions (denoted as “interp.”), yielding $\hat{s}^{G}\in\mathbb{R}^{H\times W}$. Subsequently, $\hat{s}^{G}$ is split into $n$ sub-attention maps $\{\hat{s}_{j}^{G}\}_{j=1}^{n}$ according to the local crop partitioning, where these sub-attention maps measure the semantic significance of tokens in local crops from a global perspective. Combining sub-attention maps $\{\hat{s}_{j}^{G}\}_{j=1}^{n}$ with the local importance scores $\{s_{j}^{L}\}_{j=1}^{n}$, GlobalCom2 is able to comprehensively evaluate token importance from both global and local perspectives.
Specifically, both global sub-attention maps $\{\hat{s}_{j}^{G}\}_{j=1}^{n}$ and local attention scores $\{s_{j}^{L}\}_{j=1}^{n}$ are first L2-normalized. The holistic score $s_{j,i}$ for the $i$-th token in the $j$-th local crop is then calculated by integrating its normalized global importance score $\hat{s}_{j,i}^{G}$ and local importance score $s_{j,i}^{L}$ as:
$s_{j,i}=\alpha\hat{s}_{j,i}^{G}+(1-\alpha)s_{j,i}^{L}$ (5)
Here, $\alpha\in[0,1]$ balances the two importance scores, defaulting to 0.5 for equal weighting. Based on the holistic scores, the token compression process for each local crop can be formulated as:
$\mathbf{X}^{L}_{j}\rightarrow\mathbf{\hat{X}}^{L}_{j}=\text{TopK}(\mathbf{X}^{L}_{j},s_{j},r_{j}\times N)$ (6)
where $\{\mathbf{\hat{X}}^{L}_{j}\}_{j=1}^{n}$ represents the compressed token sequences. Through holistic token evaluation, GlobalCom2 identifies tokens that are significant in global context while preserving local details, providing fine-grained visual information for LLM. See Appendix 11 for algorithm details.
4.3 GlobalCom2 without [CLS] Token
For HR-LVLMs without [CLS] token (e.g., LLaVA-OneVision [21] with SigLIP [50]), we propose an alternative token informativeness measure for GlobalCom2. Specifically, given a sequence of vision tokens $\mathbf{X}\in\mathbb{R}^{N\times d}$ after vision encoding (both thumbnail and crops), we first compute a global mean vector $\mathbf{g}\in\mathbb{R}^{d}$ through global average pooling over all tokens, and then calculate the cosine similarity between each patch token $\mathbf{x}_{i}$ and $\mathbf{g}$:
$c_{i}=\cos(\mathbf{x}_{i},\mathbf{g})=\frac{\mathbf{x}_{i}\cdot\mathbf{g}}{\|\mathbf{x}_{i}\|\|\mathbf{g}\|},$ (7)
The informativeness score $s_{i}=-c_{i}$ is negatively correlated with the similarity, reflecting that tokens with greater difference from $\mathbf{g}$ carry more unique information. Specifically, tokens exhibiting low similarity to $\mathbf{g}$ represent distinctive and irreplaceable visual elements, while highly similar tokens typically correspond to redundant or common patterns. This scoring mechanism serves as an effective alternative to [CLS]-based scoring in Eq. 2-Eq. 5 for models without [CLS] token. It is adopted to evaluate both crop-level information richness and token-level importance in LLaVA-OneVision. We conduct comprehensive quantitative and qualitative analyses in Appendix 8 to explore how to measure token informativeness without [CLS] token.
5 Experiments
Method VQAv2 GQA VizWiz SQA VQAT POPE MME MMB MMBCN MM-Vet Average
Upper Bound, 2880 Tokens
LLaVA-NeXT-7B [29] 81.8 64.2 57.6 70.1 64.9 86.5 1519.0 67.4 60.6 43.9 100.0%
Ratio=75%, Retain up to 2160 Tokens
FastV [8] 81.1 62.5 55.1 69.3 59.7 86.3 1506.3 67.6 59.0 41.7 97.5%
SparseVLM [53] 81.1 62.6 55.2 68.5 60.3 73.2 1507.8 66.1 58.6 41.9 95.7%
FasterVLM [52] 81.1 63.7 56.5 68.4 59.1 87.5 1533.4 67.5 60.2 38.5 97.4%
GlobalCom2 81.3 63.8 56.5 68.7 62.5 87.8 1548.4 68.0 60.6 40.6 98.8%
Ratio=50%, Retain up to 1440 Tokens
FastV [8] 80.7 61.8 54.9 69.1 59.6 85.5 1490.3 67.4 58.5 41.2 96.8%
SparseVLM [53] 80.9 62.0 55.7 68.1 60.0 73.4 1484.9 65.7 58.9 39.9 95.0%
FasterVLM [52] 80.6 63.4 56.4 69.1 58.9 87.7 1533.3 67.4 60.4 39.6 97.7%
GlobalCom2 80.6 63.9 56.5 68.5 62.3 88.1 1552.9 67.6 60.5 40.4 98.6%
Ratio=25%, Retain up to 720 Tokens
FastV [8] 78.9 60.4 54.2 69.8 58.4 83.1 1477.3 65.6 57.0 41.1 95.3%
SparseVLM [53] 78.9 60.9 55.6 67.5 58.1 71.0 1446.1 63.8 57.0 38.0 92.6%
FasterVLM [52] 78.3 61.3 55.4 67.1 58.8 87.2 1454.6 66.0 58.4 37.8 95.1%
GlobalCom2 79.4 61.4 55.7 68.1 60.9 87.6 1493.5 65.9 58.0 40.7 96.6%
Ratio=10%, Retain up to 288 Tokens
FastV [8] 71.9 55.9 53.1 69.3 55.7 71.7 1282.9 61.6 51.9 33.7 87.3%
SparseVLM [53] 71.6 56.1 53.2 68.6 52.0 63.2 1332.2 54.5 50.7 24.7 82.7%
FasterVLM [52] 74.0 56.9 52.6 66.5 56.5 83.6 1359.2 61.6 53.5 35.0 89.8%
GlobalCom2 76.7 57.1 54.6 68.7 58.4 83.8 1365.5 61.8 53.4 36.4 91.5%
Table 1: Comparison with other training-free LVLM token compression methods with LLaVA-NeXT across 10 benchmarks. “Average” shows the mean performance across 10 benchmarks at different retention ratios (75% to 10%), with best results highlighted.
5.1 Experimental Setting
Benchmark Details. We evaluate GlobalCom2 on various multimodal benchmarks including: VQAv2 [14], GQA [17], VizWiz [15], SQA [33], VQAT [39], AI2D [19], MMStar [7], MME [12], MMBench [32], MMBench-CN [32], MM-Vet [48], and SEED-Bench [20].
Implementation Details. We select two representative HR-LVLMs with dynamic tiling: LLaVA-NeXT [29] and LLaVA-OneVision [21] to evaluate the effects of our GlobalCom2. For hyper-parameters, $\tau$ and $\alpha$ are respectively set as 10 and 0.5 for benchmark evaluations. More implementation details are provided in the Appendix 9.
Comparison Details. We compare GlobalCom2 with FastV [8], SparseVLM [53], and FasterVLM [52] at different retention ratios $R$ (75%, 50%, 25%, 10%). While FastV and SparseVLM compress from the default 2880 tokens, FasterVLM and our method first unpad the token sequences - resulting in fewer preserved tokens given the same $R$.
5.2 Main Results
Results on LLaVA-NeXT. Table 1 compares GlobalCom2 with existing methods on 10 benchmarks using LLaVA-NeXT at different retention ratios. (1) GlobalCom2 outperforms baselines across most benchmarks, maintaining >90% of original performance at all retention ratios. Notably, GlobalCom2 even surpasses the original model on several benchmarks (MMB, MME, POPE) even at 50% retention, suggesting effective preservation of semantically crucial tokens for visual understanding. (2) GlobalCom2 excels at low retention ratios (25% and 10%), surpassing FastV and SparseVLM by 4.2% and 8.6% respectively at 10% retention. On VQAT, our method maintains stable performance while others degrade significantly (e.g., SparseVLM drops 12.9%), demonstrating its effectiveness in fine-grained visual understanding. (3) GlobalCom2 demonstrates effective hallucination mitigation on POPE, even surpassing the original model. While text-guided methods (i.e., FastV and SparseVLM) consistently show hallucination issues, vision-semantic approaches (i.e., our GlobalCom2 and FasterVLM) better preserve crucial visual details by avoiding over-reliance on instructions.
Results on LLaVA-OneVision. Figure 5 further presents GlobalCom2’s performance across benchmarks under varying retention ratios $R$ on LLaVA-OneVision. Generally, model performance correlates with $R$, with more aggressive compression leading to degradation. Vision-centric tasks (e.g., GQA, VizWiz, MME, SEED) show substantial drops with reduced visual tokens, while SQA maintains robust performance even with minimal tokens, suggesting language understanding dominates in scientific reasoning. Notably, GlobalCom2 preserves 90.5% performance at $R=10\%$ while consuming only 35.4% of the original GPU memory, all without any training overhead, demonstrating its effective and efficient token compression.
Figure 5: Results of GlobalCom2 on LLaVA-OneVision.
Figure 6: Visualization of token compression. Row 1 shows input images with thumbnail and crops. Rows 2 and 3 present compressed results from FasterVLM and our GlobalCom2 respectively, where grey masks indicate discarded tokens.
5.3 Ablation Study and Analysis
We conduct ablation studies on GlobalCom2 with $R=25\%$ across five benchmarks (VQAT, POPE, MME, SQA and MM-Vet) covering various vision-language tasks.
Ablation on Adaptive Compression Adjustment for Local Crops. Table 2 compares different compression adjustment strategies: (a) “Uniform”, a uniform strategy with fixed 25% token retention across all crops serving as our baseline, and three adaptive strategies - (b) “$\mathbf{n_{\text{top-{k}}}}$”, which adjusts compression degrees based on the proportion of thumbnail’s top-$k$ ($k=25\%\times N$) most informative tokens present in each crop region, (c) “Softmax (max)”, which applies softmax normalization over the maximum token importance score $s_{j}^{G}$ within each crop’s corresponding thumbnail region, and (d) “Softmax (sum)” (our choice), which computes softmax over the sum of token importance scores $s_{j}^{G}$ in each crop’s corresponding thumbnail region.
As shown in Table 2, our adopted “Softmax (sum)” strategy achieves the best performance. Despite “Uniform”, the other strategies acknowledge varying importance across crops. However, both $\mathbf{n_{\text{top-{k}}}}$ and “Softmax(max)” strategies only consider the strongest visual representation capability within each crop, rather than evaluating the overall importance of the entire crop to the whole image. In contrast, our adopted “Softmax (sum)” strategy dynamically modulates compression intensity according to each crop’s relative importance to the entire image, thereby preserving richer semantic visual information from high-resolution inputs and enabling LLM to better capture fine-grained visual details.
Method VQAT POPE MME SQA MM-Vet Avg.
Upper Bound, 2880 Tokens
Vanilla 64.9 86.5 1519.0 70.1 43.9 100.0%
Ratio=25%, Retain up to 720 Tokens
Uniform 60.1 87.2 1454.6 67.1 37.8 94.2%
$\mathbf{n_{\text{top-{k}}}}$ 59.8 87.3 1471.5 67.4 35.7 94.5%
Softmax (max) 60.3 87.2 1462.6 67.3 38.4 94.7%
Softmax (sum) 60.6 87.4 1473.3 67.6 39.6 95.6%
Table 2: Effects of different compression adjustment strategies.
Ablation on Holistic Token Evaluation for Local Crops. Table 3 compares different token evaluation strategies for token compression. For the $i$-th token in crop $j$, “Global” and “Local” refer to importance scores $\hat{s}_{j,i}^{G}$ and $s_{j,i}^{L}$ respectively. While both global and local token evaluation strategies are effective, each has its limitations. Local-only evaluation excels at fine-grained tasks (VQAT, POPE) but underperforms on general perception benchmarks (MME, SQA) due to missing global context. Conversely, global-only evaluation maintains good general perception but may overlook crucial local details. GlobalCom2 achieves optimal performance by combining both global and local perspectives, benefiting from their complementary strengths.
5.4 Efficiency Analysis
Table 4 compares inference efficiency among SparseVLM, FasterVLM and our GlobalCom2. Unlike SparseVLM which requires explicit attention scores during decoding and is incompatible with FlashAttention [10], FasterVLM and our GlobalCom2 enable efficient computation before LLM decoding for efficient computation. As a plug-and-play solution, GlobalCom2 achieves superior performance-efficiency trade-offs, maintaining 90% of the original performance while dramatically reducing peak memory usage by 40% and boosting inference throughput by 1.8×1.8$\times$1.8 ×. Detailed efficiency analysis can be found in Appendix 10.4.
Method VQAT POPE MME SQA MM-Vet Avg.
Upper Bound, 2880 Tokens
Vanilla 64.9 86.5 1519.0 70.1 43.9 100.0%
Ratio=25%, Retain up to 720 Tokens
Local only 60.6 87.4 1473.3 67.6 39.6 95.6%
Global only 60.2 86.4 1488.5 67.9 37.8 94.7%
Global and Local 60.9 87.6 1493.5 68.1 40.7 96.7%
Table 3: Effects of different token evaluation metrics.
5.5 Visualizations
Figure 6 visualizes token compression on VQAT, where FasterVLM [52] and our GlobalCom2 are compared under aggressive compression ($R=10\%$). FasterVLM assumes uniform information richness across crops and preserves the same ratio of tokens for each local crop, evaluating token importance only from local perspectives. This leads to both information loss and redundancy retention, as evidenced by the redundant camera metal parts preserved in the last two crops of the first case. In contrast, our GlobalCom2 adopts a global perspective to assess actual informativeness of each crop and adaptively adjusts compression intensity accordingly. This enables our method to preserve crucial visual details while maximizing redundancy removal, achieving 1.9% higher accuracy than FasterVLM on VQAT under $R=10\%$. More cases in Appendix 8 show GlobalCom2’s consistent effectiveness in adaptive token preservation across both global and local views.
Method FLOPs↓ Memory↓ Throughput↑ Performance↑
(T) (GB) (samples/s)
Upper Bound, 2880 Tokens
Vanilla 41.7 23.0 3.8 100%
Ratio=10%, Retain up to 288 Tokens
SparseVLM 5.4 (↓87.1%) 24.2 (↑5.2%) 5.9 (1.6$\times$) 82.7%
FasterVLM 3.8 (↓90.9%) 13.6 (↓40.1%) 6.7 (1.8$\times$) 89.8%
GlobalCom2 3.8 (↓90.9%) 13.9 (↓40.0%) 6.7 (1.8$\times$) 91.5%
Table 4: Efficiency comparisons. “Memory”: peak GPU memory; “Throughput”: POPE samples/second; “Performance”: average score on 10 multimodal benchmarks.
6 Conclusion
Token compression has achieved significant progress in accelerating LVLM inference. When applying existing methods to HR-LVLMs with dynamic tiling strategy, these methods treat global thumbnails and local crops uniformly, overlooking their inherent characteristics. Through analyzing HR-LVLMs with dynamic tiling, we reveal distinct roles between thumbnails and crops, and observe varying information densities across crops. Based on these findings, we propose GlobalCom2, a plug-and-play token compression framework that operates on a “global-to-local” guided principle, adaptively preserves informative regions as well as minimizes redundancy. Experiments show that GlobalCom2 achieves superior performance and efficiency, significantly outperforming existing baselines.
Supplementary Material
In the appendix, we provide theoretical FLOPs calculation in Section 7, more discussions about GlobalCom2 without [CLS] token in Section 8, detailed benchmarks introduction and implementation details in Section 9, more additional experiments and analysis in Section 10, and detailed explanation of our GlobalCom2 in Section 11.
7 Theoretical Complexity Analysis
GlobalCom2 compresses a large number of visual tokens for HR-LVLMs, thereby reducing their computational costs. Below, we analyze the theoretical computational complexity of HR-LVLMs in both the prefill stage and the decoding stage.
During the prefill stage, the FLOPs for a single transformer layer can be estimated using the formula $8Td^{2}+4T^{2}d+6Tdm$. When applying a token retention ratio $R$, where the retained token count is defined as $\hat{T}=R\cdot T$, the corresponding theoretical FLOPs reduction ratio $\eta$ can be reformulated to account for this adjustment:
$\begin{split}\eta&=1-\frac{8\hat{T}d^{2}+4\hat{T}^{2}d+6\hat{T}dm}{8Td^{2}+4T^{2}d+6Tdm}\\ &=1-\frac{R(8d+4RT+6m)}{8d+4T+6m}\end{split}$ (8)
In the decoding stage, the integration of a KV-Cache substantially enhances computational efficiency. This improvement is evidenced by the reduction in the complexity of attention computation to $\mathcal{O}(T)$. As a result, the formula for computing FLOPs is refined to $8d^{2}+4Td+6Tdm$. Given the current limitations of hardware, managing dynamic KV-Cache lengths effectively during the inference process presents significant challenges. Therefore, implementing pruning strategies prior to the decoder in large language models could facilitate a more efficient acceleration of the inference process.
8 More Discussions about Token Evaluation without [CLS].
Some Vision Encoders like SigLIP [50] do not own [CLS] tokens, rendering existing token compression methods that rely on [CLS] tokens (e.g., FasterVLM [52]) ineffective. Therefore, we explore alternative token importance evaluation approaches for such encoders. This enables GlobalCom2 to measure information content in each crop and adaptively allocate token budgets, preserving tokens with higher information richness.
As visualized in Figure 9, we first examine the attention patterns of CLIP-ViT’s final layer [CLS] token as a baseline for evaluating alternative token importance assessment strategies. We explore two approaches to measure token information without relying on [CLS] token:
(a) Patch Attention Analysis: We calculate each patch token’s importance by averaging its attention scores with all other patch tokens:
$a_{i}=\frac{1}{N-1}\sum_{j\neq i}A_{i,j},$ (9)
where $A_{i,j}$ denotes the attention score from token $i$ to token $j$, and $N$ is the total number of patch tokens. Interestingly, as shown in the “Patch Attention” visualization in Figure 9, tokens with high informativeness exhibit lower average attention scores. This aligns with the intuition that tokens heavily attending to other tokens (high average attention scores) are more likely to be replaceable, as they primarily aggregate information from their neighbors. Conversely, tokens with low average attention scores tend to be more unique and irreplaceable, containing distinct information vital for image understanding. By negating these scores (shown as “Negative Patch Attention”), we obtain patterns that relatively align with [CLS]-based assessment.
(b) Global Mean Similarity: We compute a global mean vector $\mathbf{g}\in\mathbb{R}^{d}$ through global average pooling over all tokens, then calculate the cosine similarity between each patch token $\mathbf{x}_{i}$ and $\mathbf{g}$:
$c_{i}=\cos(\mathbf{x}_{i},\mathbf{g})=\frac{\mathbf{x}_{i}\cdot\mathbf{g}}{\|\mathbf{x}_{i}\|\|\mathbf{g}\|},$ (10)
Tokens with low similarity to $\mathbf{g}$ typically represent distinctive visual elements that deviate from the average representation, suggesting their irreplaceability in capturing unique semantic information. Conversely, tokens highly similar to $\mathbf{g}$ often correspond to common or repetitive patterns, indicating their redundancy in visual representation. The “Negative Similarity with token $\mathbf{g}$” visualization aligns well with [CLS]-based assessment, highlighting semantically rich regions like the baseball player in the top-right example of Figure 9.
Based on these observations, we define two token importance measures:
$s_{i}^{attn}=-a_{i},\quad s_{i}^{sim}=-c_{i},$ (11)
where $s_{i}^{attn}$ represents the “Negative Patch Attention” score and $s_{i}^{sim}$ denotes the “Negative Similarity” score. Both scores are designed to highlight informative tokens: $s_{i}^{attn}$ identifies tokens that maintain their distinctiveness by less attending to others, while $s_{i}^{sim}$ emphasizes tokens that deviate from the average representation $\mathbf{g}$. These measures serve as effective alternatives to [CLS]-based assessment (denoted as $s_{i}^{\texttt{[CLS]}}$) in our quantitative experiments.
Method VQAT POPE MME MM-Vet MMB SQA Avg.
Upper Bound, 2880 Tokens
Vanilla 64.9 86.5 1519.0 43.9 67.4 70.1 100.0%
Ratio=25%, Retain up to 720 Tokens
$s_{i}^{\texttt{[CLS]}}$ 60.9 87.6 1493.5 39.5 65.9 68.1 96.4%
$s_{i}^{attn}$ 58.7 85.7 1449.3 36.6 63.5 67.9 93.2%
$s_{i}^{sim}$ 60.3 87.3 1485.8 39.7 65.0 67.6 95.8%
Table 5: Effects of different token evaluation metrics.
Table 5 compares the effectiveness of $s_{i}^{\texttt{[CLS]}}$, $s_{i}^{attn}$, and $s_{i}^{sim}$ when used as metrics for both adaptive compression adjustment and holistic token evaluation in GlobalCom2. While all three measures help preserve informative visual tokens to some extent, $s_{i}^{attn}$ shows notably different behavior from the other two metrics. Notably, $s_{i}^{sim}$ demonstrates performance closest to $s_{i}^{\texttt{[CLS]}}$ across various benchmarks, even outperforming it on MM-Vet. Based on these results, we adopt $s_{i}^{sim}$ as an alternative to $s_{i}^{\texttt{[CLS]}}$ for measuring information content of local crops and visual tokens when [CLS] token is unavailable, enabling training-free token compression for HR-LVLMs.
9 Detailed Experimental Settings
Benchmark Details. We evaluate GlobalCom2 on various multimodal understanding benchmarks detailed as follows:
  • GQA [17]: Contains 113,018 real-world images with scene graphs for structured visual reasoning.
  • VizWiz [15]: Features over 39,000 questions from blind users with lower-quality images and conversational style.
  • SQA [33]: Contains 21,000 scientific questions covering 26 topics and 379 different reasoning skills.
  • VQAT [39]: Features 28,000 images focusing on comprehension of text embedded in natural scenes.
  • AI2D [19]: Contains 5,000 high-resolution scientific diagrams paired with visual-spatial reasoning questions.
  • MMStar [7]: Contains 12,000 high-resolution images for spatial, temporal, and commonsense reasoning.
  • POPE [23]: Features 14,000 images with binary questions for detecting object hallucination phenomena.
  • MME [12]: Contains 15,000 images across 14 perceptual and cognitive reasoning subtasks.
  • MMB [32]: Features 20,000 multiple-choice questions designed for robust visual reasoning.
  • MMBCN [32]: Chinese counterpart of MMB with 20,000 questions for cross-lingual evaluation.
  • MM-Vet [48]: Features 16,000 questions across recognition, OCR, knowledge, and spatial reasoning.
  • SEED-Bench [20]: Contains 19,000 human-annotated questions across 12 evaluation dimensions.
Baseline Models. We select LLaVA-NeXT [29] and LLaVA-OneVision (SI) [21] as our baseline models, and follow the same inference setting as the original paper as it is publicly available [^1] LLaVA-NeXT: https://github.com/haotian-liu/LLaVA/blob/main/docs/Evaluation.md, LLaVA-OneVision: https://github.com/LLaVA-VL/LLaVA-NeXT/blob/main/docs/LLaVA_OneVision.md.. LLaVA-NeXT and LLaVA-OneVision share a common three-component architecture: a pre-trained vision encoder, a large language model (LLM) backbone, and a two-layer MLP projector bridging the two. Specifically, LLaVA-NeXT employs CLIP-ViT-L-336px [36] and Vicuna-v1.5 for vision and language modeling respectively, while LLaVA-OneVision utilizes SigLIP-So400m-Patch14-384 [50] and Qwen2 [44]. To effectively process high-resolution visual inputs, both models adopt flexible grid configurations - LLaVA-NeXT supports $\{2\times 2,1\times\{2,3,4\},\{2,3,4\}\times 1\}$ with maximum $5\times 576$ grid tokens, while LLaVA-OneVision allows $\{1\times 1,...,6\times 6\}$ grids up to $10\times 729$ tokens.
Comparison Methods. We compare our GlobalCom2 with three dominant baseline methods:
  • FastV [8] performs one-time token pruning after the selected layer of LLM based on attention map analysis.
  • SparseVLM [53] ranks token importance by text-visual attention maps, pruning through pre-selected text prompts to reduce attention noise.
  • FasterVLM [52] reranks visual tokens by [CLS] attention scores from CLIP-ViT and preserves top-$k$ tokens.
All experiments are conducted on NVIDIA A100-SXM4-80GB GPUs.
Method VQAv2 GQA VizWiz SQA VQAT POPE MME MMB MMBCN MM-Vet Average
Upper Bound, 2880 Tokens
LLaVA-NeXT-13B [29] 82.8 65.4 60.5 73.5 67.1 86.2 1575.9 70.0 64.2 48.4 100.0%
Ratio=75%, Retain up to 2160 Tokens
FasterVLM [52] 81.9 64.6 58.7 72.6 62.8 87.6 1560.3 69.8 64.8 47.5 98.7%
GlobalCom2 81.9 65.0 58.6 72.8 65.2 87.8 1567.9 69.2 64.7 49.6 99.5%
Ratio=50%, Retain up to 1440 Tokens
FasterVLM [52] 81.3 64.2 57.0 72.5 62.4 87.6 1534.1 69.5 64.1 44.7 97.3%
GlobalCom2 81.0 64.7 57.1 73.2 64.6 87.7 1553.5 69.3 63.5 48.0 98.4%
Ratio=25%, Retain up to 720 Tokens
FasterVLM [52] 78.9 62.3 55.0 72.1 61.2 86.1 1516.1 67.6 62.1 44.6 95.2%
GlobalCom2 79.9 62.7 55.1 72.3 63.6 86.5 1531.2 67.9 62.2 43.5 95.8%
Ratio=10%, Retain up to 288 Tokens
FasterVLM [52] 74.5 58.1 52.8 70.5 58.0 81.6 1386.2 61.7 53.5 38.3 88.2%
GlobalCom2 77.0 58.3 52.5 71.8 60.3 82.4 1399.5 65.0 58.5 37.9 90.3%
Table 6: Comparison with other training-free LVLM token compression methods with LLaVA-NeXT-13B across 10 benchmarks.
10 Additional Experiments
10.1 Comparisons on LLaVA-NeXT-13B.
Considering that both FasterVLM and GlobalCom2 perform token compression at the vision encoding stage, Table 6 focuses on comparing these two methods. The results reveal several key findings: (1) Our GlobalCom2 outperforms FasterVLM on most benchmarks, maintaining above 90% of uncompressed performance across different retention ratios on both LLaVA-NeXT-7B and 13B models. (2) On visual text understanding (VQAT), GlobalCom2 shows particular strength at low retention ratios, surpassing FasterVLM by 2.6% at $R=10\%$, thanks to its effective preservation of object and textual information. (3) While FasterVLM performs slightly better on general visual tasks (VQAv2, MMB, MMBCN) at high retention ratios due to its uniform token compression, GlobalCom2 demonstrates superior degradation resistance at low ratios. For example, when $R$ drops from 25% to 10%, FasterVLM’s performance significantly decreases by 5.9% and 8.6% on MMB and MMBCN, while GlobalCom2 only drops by 2.9% and 3.7%, benefiting from our carefully designed global view-guided compression strategy.
Figure 7: Hyper-parameter sensitivity analysis of $\tau$ and $\alpha$. Hyper-parameter $\tau$ controls the sharpness of probability distribution in softmax when assessing local crop informativeness, while hyper-parameter $\alpha$ determines the token retention threshold where higher values indicate stronger reliance on global guidance.
10.2 Results on LLaVA-OneVision-0.5B.
We adopt $s_{i}^{sim}$ proposed in Section 8 as the measure for crop information richness and token informativeness, and evaluate our GlobalCom2 on the LLaVA-OneVision model with more local crops across multiple benchmarks.
Our experimental results show that GlobalCom2 maintains satisfactory performance even under low retention ratios ($R=25\%,15\%,10\%$). Particularly, GlobalCom2 achieves promising results on AI2D and MMStar benchmarks which involve high-resolution images. Interestingly, we observe that on the SQA benchmark, the model performance shows minimal degradation as token retention rate decreases. We attribute this to SQA’s lower dependency on visual signals and higher reliance on LLM capabilities. Given that LLaVA-OneVision employs the powerful Qwen2 [44] as its LLM decoder, it consistently performs well on this benchmark regardless of retention ratio.
Method GQA VizWiz SQA AI2D MMStar MME SEED Avg.
Upper Bound, 7290 Tokens
LLaVA-OV 59.5 46.0 67.9 54.2 36.2 1216.1 63.8 100.0%
$R=75\%$ 58.9 44.5 67.9 52.7 37.1 1205.3 63.3 99.1%
$R=50\%$ 57.4 44.3 67.6 52.0 37.0 1216.1 62.6 98.4%
$R=25\%$ 54.2 42.1 67.0 51.3 35.8 1201.8 60.4 95.4%
$R=15\%$ 50.6 41.0 67.5 50.4 35.4 1140.2 56.8 91.1%
$R=10\%$ 48.6 40.7 67.1 50.0 35.0 1085.7 54.9 90.5%
Table 7: Results of adopting GlobalCom2 on LLaVA-OV-0.5B.
10.3 Sensitivity Analysis of Hyper-parameters.
We further explore the hyper-parameter configurations $\tau$ and $\alpha$ of our GlobalCom2 in Figure 7.
Hyper-parameter $\tau$ serves as the temperature in the softmax function when assessing local crop significance, controlling the “sharpness” of probability distribution. A smaller $\tau$ leads to a sharper distribution, amplifying the differences in global visual importance among local crops. As shown in the first row of Figure 7, our GlobalCom2 demonstrates robust performance across most benchmarks, particularly on VQAT [39] and POPE [23], with varying $\tau$ validating our design principle for HR-LVLMs of allocating retention ratios based on each local crop’s global importance. We empirically set $\tau=10$ to achieve optimal performance across most benchmarks.
Hyper-parameter $\alpha$ determines the criterion for token retention in local crops, where a larger $\alpha$ indicates stronger dependence on global guidance for token retention in local crops. Similar to $\tau$, the second row of Figure 7 shows that our GlobalCom2 exhibits consistent effectiveness across most benchmarks under different $\alpha$ settings, substantiating our dual-perspective token retention strategy of local crops by considering their significance at both global and local levels for HR-LVLMs.
Method TFLOPs↓ Peak Memory (GB)↓ KV-Cache (MB)↓ Prefill Time (ms)↓ Throughput (samples/s)↑ Performance↑
Upper Bound, 2880 Tokens
LLaVA-NeXT-7B 41.7 23.8 1536.0 170.7 2.5 1519.0
Ratio=75%, Retain up to 2160 Tokens
GlobalCom2 30.4 (↓27%) 17.8 (↓25%) 1126.4 (↓27%) 119.9 (↓30%) 3.2 (1.3$\times$) 1548.4
Ratio=50%, Retain up to 1440 Tokens
GlobalCom2 19.7 (↓53%) 16.2 (↓32%) 755.0 (↓51%) 74.5 (↓57%) 4.2 (1.7$\times$) 1552.9
Ratio=25%, Retain up to 720 Tokens
GlobalCom2 9.6 (↓77%) 14.8 (↓39%) 377.0 (↓76%) 34.6 (↓80%) 5.3 (2.1$\times$) 1493.5
Ratio=10%, Retain up to 288 Tokens
GlobalCom2 3.8 (↓91%) 14.2 (↓40%) 151.0 (↓90%) 13.3 (↓92%) 6.8 (2.7$\times$) 1365.5
Table 8: Efficiency analysis of our GlobalCom2 with LLaVA-NeXT-7B on one NVIDIA A100-SXM4-80GB GPU.
Method TFLOPs↓ Peak Memory (GB)↓ KV-Cache (MB)↓ Prefill Time (ms)↓ Throughput (samples/s)↑ Performance↑
Upper Bound, 2880 Tokens
LLaVA-NeXT-13B 80.0 36.7 2457.6 295.1 1.9 1575.9
Ratio=75%, Retain up to 2160 Tokens
GlobalCom2 58.7 (↓27%) 34.1 (↓7%) 1843.2 (↓25%) 211.6 (↓28%) 2.4 (1.3$\times$) 1567.9
Ratio=50%, Retain up to 1440 Tokens
GlobalCom2 38.3 (↓52%) 30.1 (↓18%) 1228.8 (↓50%) 134.6 (↓54%) 3.2 (1.7$\times$) 1553.5
Ratio=25%, Retain up to 720 Tokens
GlobalCom2 18.7 (↓77%) 27.5 (↓25%) 590.0 (↓76%) 64.1 (↓78%) 4.8 (2.4$\times$) 1531.2
Ratio=10%, Retain up to 288 Tokens
GlobalCom2 7.4 (↓91%) 26.2 (↓29%) 236.0 (↓90%) 25.0 (↓92%) 5.8 (3.1$\times$) 1399.5
Table 9: Efficiency analysis of our GlobalCom2 with LLaVA-NeXT-13B on one NVIDIA A100-SXM4-80GB GPU.
10.4 Detailed Efficiency Analysis.
We conduct a comprehensive analysis of both theoretical and practical efficiency of our GlobalCom2 with LLaVA-NeXT-7B/13B on a single NVIDIA A100-SXM4-80GB GPU, as shown in Table 8 and Table 9. We measure computational efficiency using TFLOPs and peak memory consumption directly, while other metrics are estimated using LLM-Viewer [49]. Given that the sequence length of visual tokens substantially exceeds that of textual and system tokens, we exclude the latter two from our theoretical analysis. For practical efficiency, “Throughout” and the corresponding “Performance” measurements are conducted on the MME benchmark [12], which includes 2374 examples.
As shown in Table 8 and Table 9, GlobalCom2 substantially improves the computational efficiency of LLaVA-NeXT models across both theoretical and practical metrics. Specifically, while preserving model performance, our method achieves significant reductions in GPU memory consumption and notable acceleration in inference speed compared to the original LLaVA-NeXT. Most importantly, these efficiency improvements are achieved through a training-free compression scheme, demonstrating the immediate practical value of GlobalCom2.
10.5 More Visualizations of Token Compression
Figure 8: More visualization of token compression by GlobalCom2. The presented examples are from VQAT, where grey masks indicate discarded visual tokens.
In Figure 8, we present more token compression visualizations of GlobalCom2. It is clear that in all cases, entity-rich regions are effectively preserved while redundant ones are removed. For crops with high visual redundancy (e.g., the camera’s golden casing in the top-right case), it applies aggressive compression ($r_{4}=19\%$). Conversely, for semantically dense regions (e.g., the text-rich first crop in the top-right case), it maintains more tokens ($r_{1}=43\%$) to ensure detailed understanding. More cases in Appendix 8 further demonstrate GlobalCom2’s consistent ability to adaptively preserve significant regions while removing redundancy at both global and local views.
11 Algorithm Illustration
We present a comprehensive description of GlobalCom2 token compression for both global thumbnail and local crops in Algorithm 1 and Algorithm 2, respectively.
Figure 9: Visualization of different token evalutaion scores.
Algorithm 1 GlobalCom2: Thumbnail Compression
1:Global thumbnail tokens $\mathbf{X}^{G}\in\mathbb{R}^{N\times D}$, Vision encoder, Preset retention ratio $R\in(0,1]$
2:Compressed tokens $\mathbf{\hat{X}}^{G}\in\mathbb{R}^{k\times D}$ ($k=R\cdot N$)
3:// For models with [CLS] token
4:if model has [CLS] token then
5:       Get [CLS] query $\mathbf{q}^{\texttt{CLS}}$ and Key matrix $\mathbf{K}$
6:       Compute importance scores:
$s_{i}^{G}\leftarrow\frac{\exp(\mathbf{q}^{\texttt{CLS}}\mathbf{K}_{i}^{\top}/\sqrt{D})}{\sum_{j=1}^{N}\exp(\mathbf{q}^{\texttt{CLS}}\mathbf{K}_{j}^{\top}/\sqrt{D})},\forall i$
7:else
8:       // For models without [CLS] token
9:       Compute global mean vector: $\mathbf{g}\leftarrow\frac{1}{N}\sum_{i=1}^{N}\mathbf{x}_{i}$
10:       Compute cosine similarity scores:
$s_{i}^{G}\leftarrow-\frac{\mathbf{x}_{i}\cdot\mathbf{g}}{\|\mathbf{x}_{i}\|\|\mathbf{g}\|},\forall i$
11:end if
12:Sort indices: $\text{idx}\leftarrow\text{argsort}([s_{1}^{G},...,s_{N}^{G}])$ (descending)
13:Retain top-$k$ tokens: $\mathbf{\hat{X}}^{G}\leftarrow\mathbf{X}^{G}[\text{idx}[1:k]]$
14:return $\mathbf{\hat{X}}^{G}$
Algorithm 2 GlobalCom2: Local Crop Compression
1:Local crops $\{\mathbf{X}_{j}^{L}\}_{j=1}^{n}$, Global importance scores $s^{G}$, Preset retention ratio $R$, Balance weight $\alpha$, Temperature $\tau$
2:Compressed crop tokens $\{\mathbf{\hat{X}}_{j}^{L}\}_{j=1}^{n}$
3:Adaptive Compression Adjustment:
4:for each crop $j\in[1,n]$ do
5:       Compute information richness:
$s_{j}^{G}\leftarrow\sum_{i\in\text{crop}_{j}}s_{i}^{G}$
6:       Normalize: $\tilde{s}_{j}\leftarrow(s_{j}^{G}-\max\{s_{j}^{G}\})/\tau$
7:end for
8:Compute importance weights:
$\sigma_{j}\leftarrow\frac{\exp(\tilde{s}_{j})}{\sum_{l=1}^{n}\exp(\tilde{s}_{l})+\epsilon}$
9:Adjust retention ratios:
$r_{j}\leftarrow R\times(1+\sigma_{j}-\frac{1}{n})$
10:Holistic Token Evaluation:
11:for each crop $j\in[1,n]$ do
12:       Compute local scores $s_{j}^{L}$ via vision encoder
13:       Get global sub-scores $\hat{s}_{j}^{G}$ via interpolation
14:       L2-normalize both $s_{j}^{L}$ and $\hat{s}_{j}^{G}$
15:       Combine scores: $s_{j,i}\leftarrow\alpha\hat{s}_{j,i}^{G}+(1-\alpha)s_{j,i}^{L}$
16:       Compress: $\mathbf{\hat{X}}_{j}^{L}\leftarrow\text{TopK}(\mathbf{X}_{j}^{L},s_{j},r_{j}\times N)$
17:end for
18:return $\{\mathbf{\hat{X}}_{j}^{L}\}_{j=1}^{n}$